OpenStudy (aamirgsen):

The measure of the complement of angle A is 185 less than two times the measure of the supplement of angle A. Find measure of angle A. All help appreciated :D

OpenStudy (bibby):

complementary angles => A+B=90 complementary angles => A+B=180 the complement of angle A=(90-A) the supplement of angle A=(180-A)